# LLM-3D Print: Large Language Models To Monitor and Control 3D Printing

### v1

@@ 4.2 Error Detection
The participants were asked to identify major issues in the ongoing prints, such as layer adhesion, surface quality, and defects like under-extrusion, stringing, or warping. Their evaluations were then compared to the LLM’s automated detection and optimization process to determine how effectively the LLM identified and addressed these issues in real-time. This comparison allowed us to gauge not only the LLM’s accuracy in detecting and mitigating errors but also how its performance measured up against human expertise at various levels of experience.
To quantify the effectiveness of the LLM’s error detection relative to human evaluations, we introduce the metrics of agreement rate and disagreement rate. These metrics are defined as follows:
$AgreementRate=\sum_{i}^{{N}}\frac{{D_{P}}}{{Total_{P}}}*{D_{LLM}}$
$DisagreementRate=-\sum_{i}^{{N}}\frac{{D_{P}}}{{Total_{P}}}*{\hat{D}_{LLM}}$
Here, $N$ represents the total number of layers in the print, $D_{P}$ denotes the number of participants who identified a particular defect, and $Total_{P}$ is the total number of participants. $D_{LLM}$ indicates whether the LLM detected the defect (initialized to 0 and set to 1 if the LLM identifies a defect), while $\hat{D}_{LLM}$ represents defects not detected by the LLM (initialized to 1 and set to 0 if the LLM identifies the defect).
The agreement rate measures the alignment between the defects identified by the LLM and those recognized by the participants, reflecting the LLM’s accuracy in matching human observations. Conversely, the disagreement rate captures instances where the LLM failed to detect defects noted by participants, highlighting areas where the LLM’s performance may need improvement. These metrics provide a comprehensive assessment of the LLM’s ability to identify and correct printing defects compared to human evaluations.
Figure 6 b presents a comparison between participant responses and the LLM’s detection results. The analysis shows that both the LLM and the majority of participants consistently identified stringing (77%) and under-extrusion (55%) as significant issues across most print layers. Additionally, the LLM accurately detected more critical defects such as bed adhesion problems, warping, and layer separation.
The LLM did not detect some of the more subtle defects, such as print cracks, ghosting, ringing, and elephant foot, which were noted by approximately 1% of the participants. This discrepancy could be due to the resolution limitations of the camera used for image capture, which may not have been sufficient to detect these finer details. Another possibility is that these minor issues were misidentified by the participant evaluators, resulting in a difference between human and LLM detection.
Interestingly, the LLM successfully identified blobs and zits in layers 10 and 15, but it attributed these issues to stringing and oozing defects in other layers. This suggests that while the LLM is effective at detecting and categorizing certain types of defects, it may occasionally generalize or misclassify closely related issues, particularly when dealing with subtle variations in print quality. This highlights both the LLM’s strengths in detecting major defects and areas where further refinement is needed to improve its sensitivity to minor or less common issues.
Based on the agreement-disagreement scores, it can be observed that the LLM and human evaluators show a 77% agreement on identifying stringing/oozing defects, a 20% agreement on under-extrusion, and a 7% agreement on over-extrusion. Additionally, there is a maximum of 20% disagreement between the LLM and humans for blobs and zits defect. These figures indicate that the LLM demonstrates strong alignment with human assessments for key visual defects like stringing/oozing and inconsistent extrusion. The close agreement on critical defects shows the LLM’s potential, with room for growth in more nuanced areas.

### v2

@@ 5.2 LLM Identified Defects
To further evaluate the effectiveness of the LLM (GPT-4o) in detecting errors during multi-layer prints, we conducted a user study involving a diverse cohort of 14 engineers, with varying levels of experience in additive manufacturing. Participants ranged from individuals with basic exposure to AM to those with advanced domain expertise. Each participant was asked to evaluate multi-layer prints that had been previously analyzed and optimized by the LLM. Participants were asked to identify major issues in the ongoing prints, such as poor layer adhesion, surface artifacts, and common defects like under-extrusion, stringing, and warping. Their evaluations were then compared to the LLM’s automated detection outputs to assess how effectively the model identified and identified these issues. This comparison enabled us to evaluate not only the LLM’s accuracy in detecting and mitigating print defects, but also how its performance compared to human expertise across varying levels of experience.
Figure 7b presents a comparative analysis of participant responses, expert annotations, and the LLM’s detection results. The analysis shows that both the LLM and the majority of participants, including experts, consistently identified stringing and extrusion-issues as prominent failure modes across multiple layers. In addition, the LLM was able to detect more complex defects, such as bed adhesion failures, warping, and layer separation, which were sometimes overlooked by less experienced participants. These results suggest that the LLM not only matches near-human-level diagnostic capabilities in common error scenarios but can also offer enhanced detection in more subtle or cumulative failure conditions, reinforcing its role as a reliable agent in automated quality monitoring.
It is important to note that some print defects such as cracks, ghosting, ringing, and elephant foot, were reported by a small subset of participants but were not confirmed by expert annotations. This discrepancy likely reflects either over-attribution by less experienced participants or the inherent subjectivity involved in assessing subtle visual anomalies. The relatively low frequency of these responses further suggests that such defects were either inconsistently perceived or misclassified. This divergence between participant feedback and expert assessments reinforces the importance of using expert-labeled annotations as a consistent and reliable ground truth for evaluating the performance of automated systems like the LLM.
Figure 8: Confusion matrix. Confusion matrices comparing detected print defects against expert annotations. (Left) LLM-predicted defects show high true positive rates for major issues such as extrusion errors and stringing/oozing, but also include a few false positives and misclassifications, particularly for visually similar anomalies like blobs/zits. (Right) Human survey responses show greater variability, with more false positives in subtle defect categories such as print cracks, ghosting, and elephant foot—indicating potential over-identification or misclassification by less experienced participants.
Interestingly, the LLM correctly identified blobs and zits in layers 10 and 14, yet it attributed similar patterns in other layers to stringing or oozing. This suggests that while the LLM demonstrates strong capabilities in recognizing and categorizing major defects, it may sometimes generalize or conflate closely related print anomalies. Such behavior points to both the promise and current limitations of LLMs in fine-grained defect classification, highlighting the need for further refinement to enhance sensitivity to minor or less common print issues.
In some cases, the LLM outperformed human evaluators. For example, it detected early signs of layer separation in layers 7 and 11 before they were identified by expert annotators. This indicates that the model can sometimes anticipate structural failures earlier than humans by reasoning over visual progression across layers.
However, certain defects such as minor warping were occasionally misattributed, and subtle anomalies like layer separation were not always consistently labeled. This may be partly due to limitations in image resolution, as the dataset was compressed to remain compatible with LLM token constraints. Consequently, defects that require high visual fidelity may not be fully captured in the input data, restricting the model’s ability to identify them with high accuracy.
A more quantitative comparison of detection performance is provided in Figure 8, which presents confusion matrices for both the LLM and human survey responses against expert annotations. The LLM demonstrates strong precision and recall in identifying high-impact defects such as extrusion issues and stringing/oozing, with few false positives or false negatives. In contrast, human survey responses showed greater variability, particularly in the classification of subtle defects such as ghosting, elephant foot, and print cracks. Participants frequently over-identified these rare defects, leading to a higher rate of false positives. This further highlights the LLM’s ability to maintain diagnostic accuracy while minimizing over-attribution.
5.3 Generalization Without Supervision
The demonstrated ability to detect and correct 3D printing defects highlights the potential of our approach as a general-purpose solution for process monitoring in additive manufacturing. Unlike conventional deep learning methods that require large annotated datasets and task-specific retraining, our framework leverages prompt-based reasoning and in-context learning to operate without any fine-tuning. The LLM interprets visual cues, queries printer state, and generates corrective actions directly from multimodal inputs. While our evaluation was conducted on a per-layer sampling basis, the framework can be readily extended to operate over multi-layer intervals to reduce the frequency of LLM invocations. Currently, processing a single layer takes between 15 to 45 seconds, depending on the complexity of the failure (judged by number of tokens generated) and whether the model successfully identifies the relevant API endpoints. This latency can be further reduced by explicitly providing the LLM with structured endpoint mappings for each sensor or parameter, improving both response time and reliability.
Recent efforts in automated defect detection have primarily focused on training deep learning models with large-scale annotated datasets, yet these efforts often optimize only for a narrow range of simple failure modes. For instance, one approach trained a convolutional neural network (CNN) on over 120,000 labeled images, using only three coarse defect categories: “Good-quality,” “Under-extrusion,” and “Over-extrusion” 38. Similarly, a vision transformer (ViT)-based model relied on heavily curated data and extensive training to identify a limited set of extrusion defects 106. While such models perform well within their labeled domains, their effectiveness is bounded by the specificity of the annotations. They struggle to generalize to new failure types, printer hardware, or parameter variations not explicitly present in the training data. While the LLM may occasionally misattribute certain defects on a per-layer basis, such as confusing blobs with stringing or overgeneralizing failure categories, its aggregate performance across entire prints leads to significantly improved structural outcomes. By consistently identifying key defect patterns and applying corrective parameter adjustments, the LLM contributes to enhanced print quality over time.
In contrast, our LLM-based approach leverages general reasoning capabilities and sequential prompting to detect and diagnose a much broader set of failure modes, including stringing, layer separation, warping, and bed adhesion issues. It attributes faults to relevant printer parameters and proposes corrective actions, without relying on domain-specific supervision or pre-labeled examples. This transition from supervised learning to inference-driven reasoning marks a fundamental shift in approach, enabling scalable, interpretable, and adaptable quality control. By removing the dependence on annotated datasets and retraining cycles, our method represents a more flexible and cost-effective pathway for deploying intelligent monitoring in additive manufacturing environments.
A key advantage of our framework is its modular design, which allows seamless integration of different error detection modules without requiring substantial changes to the overall system. While we demonstrate the utility of a LLM for detection, the architecture can accommodate alternative models, such as domain-specific vision networks, heuristic algorithms, or hybrid sensor fusion modules, simply by replacing the error detection component or any other relevant component. This modularity not only promotes extensibility but also future-proofs the system against advances in detection algorithms.
Our framework was deployed on a standard consumer-grade FDM printer with a 1 mm nozzle, lacking manufacturer-provided settings. The LLM successfully diagnosed and corrected print issues without device-specific tuning. Its reasoning capabilities extended across incomplete inputs and generalized settings. The framework also adapts dynamically to available sensor data, including temperature, retraction settings, print speed, and G-code offsets, based on firmware accessibility. This sensor-agnostic design ensures compatibility with diverse printer models and configurations, supporting wide applicability without requiring hardware-specific integration.